The sum of the ages of Abu and Martin is 30 years.Abu is 6 years younger than Martin.How old is Martin?
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

The sum of the ages of Abu and Martin is 30 years.Abu is 6 years younger than Martin.How old is Martin?
===============
M + A = 30
M - A = 6
------------------Add

Answer by ikleyn(52834)

.

Martin + Abu        = 30


Martin + (Martin-6) = 30


2*Martin = 30 + 6 = 36

 
  Martin          = 36/2 = 18 years old.    ANSWER